OpenStudy (aonz):

Log question. Is there a rule that can solve this? \[\huge \log_{\frac{ 1 }{ a }} \frac{ 1 }{ n }\]

OpenStudy (aonz):

@agent0smith can you help with this?

OpenStudy (agent0smith):

Yeah i'm not sure what you want to do with it... you can change it into a couple of different forms?

OpenStudy (anonymous):

There is a formula for changing the base of the logarithm.

OpenStudy (aonz):

well il type full equation prove log a(x) = -log 1/2 (x)

hartnn (hartnn):

use : \(\huge \log_ab=\dfrac{\log b}{\log a}\)

hartnn (hartnn):

u sure, this is the Q ? prove log a(x) = -log 1/2 (x)

OpenStudy (aonz):

wait my bad LOL log a(x) = -log 1/a (x)

hartnn (hartnn):

right, use the property i gave on RIGHT side.

OpenStudy (aonz):